Find the general term in the expansion of (2-3x)-1/2?

Grade:12

3 Answers

OUTPASS
8 years ago
Given (2-3x)-1/3
(2-3x)-1/3   = 2-1/3(1-3x/2)-1/3
finally;.the answer is 2-1/31*4*7*........*(3r-2)/r|
venky
8 years ago
 
Given (2-3x)-1/3
(2-3x)-1/3   = 2-1/3(1-3x/2)-1/3
finally;.the answer is 2-1/31*4*7*........*(3r-2)/r|
RAJU
8 years ago
(2-3x)-1/2
=2-1/3(1-3x/2)-1/3
Tr+1=2-1/3 [{(-1/3)(-1/3-1)(-1/3-2)...........(-1/3-r+1)}/r!](-3x/2)r
=(2-1/3/r!)(-12r)(1/3)(4/3)(7/3).................(3x/2)r
=(2) (1*4*7*...............*(3r-2)/r!)(x/2)r